OpenStudy (anonymous):

i just need to solve 0.037=x+1 please help

OpenStudy (anonymous):

i think i need to divide both sides by x+1

Vocaloid (vocaloid):

Subtract 1 from each side

OpenStudy (anonymous):

-0.963?

OpenStudy (anonymous):

@Vocaloid could you help me with 2 more?

Vocaloid (vocaloid):

Sure

OpenStudy (anonymous):

The Mendes family bought a new house 9 years ago for $108,000. The house is now worth $165,000. Assuming a steady rate of growth, what was the yearly rate of appreciation? Round your answer to the nearest tenth of a percent (1.2% etc)

Vocaloid (vocaloid):

First find (change in price)/(number of years)

OpenStudy (anonymous):

18333.33

Vocaloid (vocaloid):

(165000-108000)/9 = ?

OpenStudy (anonymous):

6333.33

Vocaloid (vocaloid):

right, now we find the percentage compared to the original price 100*(6333.33/108000) = ?

OpenStudy (anonymous):

5.864?

Vocaloid (vocaloid):

right, then round to the nearest 10th

OpenStudy (anonymous):

5.9?

Vocaloid (vocaloid):

right, so our answer is 5.9%

OpenStudy (anonymous):

An anthropologist finds bone that her instruments measure it as 0.494% of the amount of Carbon-14 the bones would have contained when the person was alive. How long ago did the person die? The half life of carbon 14 is 5,730 years. Round your answer to the nearest thousand.

Vocaloid (vocaloid):

portion remaining = (2)^(number of years/half-life) .494/100 = (2)^(x/5730) solve for x

OpenStudy (anonymous):

i couldnt solve it

OpenStudy (anonymous):

@Vocaloid help, i couldnt solve it

Vocaloid (vocaloid):

.494/100 = (2)^(x/5730) log(.494/100)/log(2) = x/5730

OpenStudy (anonymous):

@Vocaloid keep getting error and i cant solve it :(

Vocaloid (vocaloid):

small correction log(.494/100)/log(0.5) = x/5730 7.66 = x/5730 x = 44,000 years (rounded)

OpenStudy (anonymous):

okay one more easy one e^2x+1=5.324

OpenStudy (andrearchat):

X=-.73

OpenStudy (anonymous):

1.672=2x+1 , now what? @Vocaloid

OpenStudy (anonymous):

wait no its 1.672=2x-1

Vocaloid (vocaloid):

 is this what the problem looks like?